Prove that every non-trivial subgroup of a cyclic group has finite index. Hence 

prove that (Q, +) is not cyclic.


1
Expert's answer
2021-06-07T05:03:27-0400

Let us prove that every non-trivial subgroup of a cyclic group has finite index. If a cyclic group is finite, then all its subgroup finite and have finite index. If a cyclic group is infinite, then it is isomorphic to the additive group of integers. Taking into account that all subgroups of a cyclic group are cyclic, we conclude that every non-trivial subgroup of a cyclic group is of the form "k\mathbb Z" for "k\in\mathbb Z\setminus\{0\}." The quotient group "\mathbb Z/k\mathbb Z=\{[0],[1],\ldots [k-1]\}" is of order "k", and hence every non-trivial subgroup of a cyclic group has finite index.


The group "(\mathbb Q, +)" contains non-trivial subgroup "(\mathbb Z, +)" of infinite continuum index: "|\mathbb Q/\mathbb Z|=|[0,1)|=c." Hence the group "(\mathbb Q, +)" is not cyclic.